OpenStudy (anonymous):

40 is what percent of 50?

OpenStudy (anonymous):

100

OpenStudy (anonymous):

sorry its 125

OpenStudy (doc.brown):

It is not 100 or 125 @brown_562

OpenStudy (doc.brown):

\[\dfrac{40}{50}=\dfrac{0.8}{1}\]You want how many it is out of 100.\[\dfrac{0.8}{1}\times\dfrac{100}{100}=\dfrac{80}{100}\]

OpenStudy (cggurumanjunath):



OpenStudy (anonymous):

80 % it is.